OpenStudy (anonymous):

what is the solution to this system of linear equations? y=-6x-47 y=2x+17

OpenStudy (hoblos):

-6x-47=2x+17 -8x = 64 x =-8 y=2x+17 = -16+17 = 1

OpenStudy (anonymous):

y=9x+32 y=-9x-40

OpenStudy (hoblos):

9x+32 = -9x-40 18x = -72 x = -4 y=9x+32 = -36+32 = -4

OpenStudy (anonymous):

y=-2x+6 y=3x+6

OpenStudy (anonymous):

i try to do it but i dont fully understand how to

OpenStudy (hoblos):

-2x+6 = 3x+6 subtract (3x) from both sides -5x +6 = 6 subtract (6) from both sides -5x=0 divide by (-5) x=0 y=-2x+6 replace (x=0) y=6

OpenStudy (hoblos):

does this make it clear to you

OpenStudy (anonymous):

yes thank you

OpenStudy (hoblos):

my pleasure XD